OpenStudy (anonymous):

–5 + 2x² = –6x ? helpppppp medal given solve problem

OpenStudy (anonymous):

2x^2+6x-5=0 x=(-6+-sqrt(6^2-4*2*-5)/2*2 =(-6+-sqrt76)/4 =(-6+-2sqrt19)/4 =(-3+-2sqrt19)/2 now you can solve.

OpenStudy (anonymous):

solve what what do you mean the answer please ?

OpenStudy (haseeb96):

2x^2 +6x -5=0

OpenStudy (anonymous):

x=(-3+2sqrt19)/2 and x=(-3-2sqrt19)/2

OpenStudy (haseeb96):

here u can apply Quadratic formula

OpenStudy (anonymous):

if you want to solve it by making factors then let me know.

OpenStudy (haseeb96):

she can't be made factors because the middle term will not be broken ...

OpenStudy (anonymous):

x^2+3x-5/2=0 or x^2+3x+(3/2)^2-5/2-(3/2)^2=0 or (x+3/2)^2-5/2-9/4=0 or (x+3/2)^2-19/4=0 or (x+3/2)^2-{(sqrt19)/2}^2=0 or {x+3/2+(sqrt19)/2}{x-3/2-(sqrt19)/2}=0 now you can find the value of x.